OpenStudy (anonymous):

Given a point (2, 3) and the slope of a line M = 4, find the equation of the line.

OpenStudy (anonymous):

y = 4x + c now as the line passes through point (2, 3), it must satisfy the equation so 3 = 4 (2) + c c = 3 - 8 c = -5 so required equation is y = 4x -5

OpenStudy (anonymous):

thank you :)